Using the Epsilon and Delta definition, show that lim x approaching 2 for (3x-5) is =1



1
Expert's answer
2021-11-07T18:08:24-0500

"Given \ any \ \epsilon \gt0 \ we \ choose\ \delta_\epsilon \lt \frac{\epsilon}{3}"

Explanation

we have to show or prove that given any "\epsilon \gt0" we can find a "\delta_\epsilon \gt 0" such that:

"\lvert f(x) -1 \rvert \lt \epsilon \space \ for \ x \isin (2- \delta_\epsilon , 2+ \delta_\epsilon )"


we evaluate the difference:

"\lvert f(x) -1 \rvert = |3x-5-1|=|3x-6|\\=3|x-2|"


and we can see that :


"|f(x)-1|\lt \epsilon \iff |x-2| \lt \frac{\epsilon}{3}"


so given "\epsilon \gt 0" we can choose "\delta_\epsilon \lt \frac{\epsilon}{3}" and we have:

"x \isin (2- \delta_\epsilon , 2+ \delta_\epsilon ) \implies |x-2| \lt \frac{\epsilon}{3}\\\iff |f(x)-1|\lt\epsilon"


which shows the point.